# myboxenhanced,colback=blue!5!white,colframe=blue!75!black,boxrule=0.6mm,fontupper=

### v1

@@ 6.2 Unsupervised RL
Structural decompositions can help such methods by improving the pre-training phase’s tractability while improving the fine-tuning phase’s generality. Latent decompositions could help exploit structure in unlabeled data, while relational decompositions could add interpretability to the learned representations. Through augmentation, conditioning policies on specific parts of the state space can reduce the amount of data needed for fine-tuning. Additionally, understanding problem decomposition can simplify complex problems into more manageable sub-problems, effectively reducing the perceived problem complexity while allowing the incorporation of such decomposition in external curricula for fine-tuning. Incorporating warehousing guided by decompositions for competence-based methods can boost the fine-tuning process of the learned skills.
{mybox}
Unsupervised RL Patterns
  • Use latent decompositions to extract structure from unlabeled data, reducing Data Dependency.
  • Employ factored and modular decompositions and abstractions to manage scalability by focusing learning on different parts of the problem independently.
  • Warehouse skills across different modular sub-problems to reuse solutions and enhance Generality.
  • Manage Problem Complexity by leveraging problem decomposition to simplify the learning task and using decompositions for fine-tuning using curriculum learning.
@@ 6.3 Big Data and Foundation models in RL
Foundation models for RL come increasingly closer to becoming a reality. Such RL models would follow a similar concept of training a large model on a wide variety of tasks, environments, and behaviors to be fine-tuned for specific downstream tasks. SMART (?), one of the current contenders for such models, follows this paradigm by using a self-supervised and control-centric objective that encourages the transformer-based model to capture control-relevant representation and demonstrates superior performance when used for fine-tuning. AdA (?) trains an in-context learning agent on a vast distribution of tasks where the task factors are generated from a latent ruleset.
Given the pre-training paradigm, these methods are highly data-dependent in principle. However, incorporating large amounts of data can demonstrate scalability benefits by reducing fine-tuning costs for distributed applications. A natural question that arises is the role of Structured RL in the realm of end-end learning and big data. Even though such methods subscribe to an end-to-end paradigm, structural decompositions can benefit them in different ways.
@@
Researchers can better understand the decomposability in the pre-trained models by categorizing methods based on how they incorporate structure. Consequently, this can guide the selection processes for fine-tuning methods depending on the tasks at hand. Passive learning from pre-trained models can benefit from better explanations about what parts of a fine-tuning task space might be suited for what kind of warmstarting strategies. Additionally, incorporating interpretability-oriented decompositions such as relational representations can help design more interpretable fine-tuning methods.
@@
Structural information can guide the development of novel architectures. With a better understanding of how different architectures and algorithms incorporate structural information, practitioners can more effectively adapt existing methods or contribute to designing novel solutions tailored to their specific tasks. For example, Action Branching architectures might provide modular functionality in downstream tasks, especially suited for multi-task settings. On the other hand, representation bottlenecks might suit settings that deviate from each other by small changes in contextual features.
@@
By understanding how to decompose tasks and incorporate structural information, foundational models can be fine-tuned more effectively for specific tasks or transferred to new tasks. The understanding of decompositions could guide how to structure the fine-tuning process or how to adapt the foundational model to a new task. By understanding how to incorporate structural information during fine-tuning, they can potentially achieve improved performance.
@@
By understanding the spectrum of decomposability and how various methods incorporate structure, we can create better benchmarks and evaluation protocols for foundational models. For instance, we can evaluate how well foundational models handle tasks of different decompositions and patterns of fine-tuning. Researchers can use this framework to design better evaluation protocols and benchmarks for foundational models. For practitioners, such benchmarks and evaluation protocols can guide the selection of models and algorithms for their specific tasks.
{mybox}
Foundation Model Patterns
  • Use Factored or Relational abstractions on the pre-trained foundation model for state abstractions to manage high-dimensional state spaces and, thus, reduce data dependency.
  • Condition the policy on additional task-specific information, such as goal information, representation of specific fine-tuning instructions, or control priors to improve scalability.
  • Regularize the fine-tuning process to prevent catastrophic forgetting of useful features learned during pre-training.
  • Maintain a warehouse of fine-tuned policies and value functions to help reuse previously learned skills and adapt them to new tasks, improving learning efficiency and generalization.
  • Incorporate a curriculum of increasingly complex fine-tuning environments based on the agent’s performance to help the agent gradually adapt the foundation model’s knowledge to the specific RL task.
  • Use explicit architectures that fine-tune different RL problem aspects, such as perception, policy learning, and value estimation.
@@ 6.4 Partial observability and Big Worlds
In many real-world situations, the Markov property might not fully capture the dynamics of the environment (?, ?). This can happen in cases where the environment’s state or the rewards depend on more than just the most recent state and action or if the agent cannot fully observe the state of the environment at each time step. In such situations, methods need to deal with non-Markovian dynamics and partial observability.
@@
Abstractions can play a crucial role in such situations, where structural decompositions using abstraction patterns can make methods more sample efficient. Often used in options, temporal Abstractions allow the agent to make decisions over extended periods, thereby encapsulating potential temporal dependencies within these extended actions. This can effectively convert a non-Markovian problem into a Markovian one at the level of options. State Abstractions abstract away irrelevant aspects of the state and, thus, can sometimes ignore certain temporal dependencies, rendering the process Markovian at the level of the abstracted states. Thus, research into the role of decompositions in abstraction opens up possibilities to understand the dependencies between non-Markovian models and the kind of abstractions they use to solve problems with incomplete information. Abstraction can also simplify the observation space in POMDPs, reducing the complexity of the belief update process. The abstraction might involve grouping similar observations together, identifying higher-level features, or other simplifications. Abstractions can additionally allow us to break partial observability down into different types instead of always assuming the worst-case scenario. Utilizing such restricted assumptions on partial observability can help us build more specific algorithms and derive convergence and optimality guarantees for such scenarios.
@@
Any additional information required, such as belief states or memory of past observations, can be used as abstractions or augmentations. This can also help with more efficient learning of transition models for planning. Hierarchical techniques that utilize optimization at different timescales can incorporate warehousing to reuse learned policies across various levels of abstraction. Environment generation patterns could also be used to generate a curriculum of increasingly complex tasks for the agent, starting with simpler MDPs and gradually introducing partial observability or other non-Markovian features.
@@
As we extend the information content of the environment to its extremity, we delve into the realm of the big world hypothesis in RL (?), where the agent’s environment is multiple orders of magnitude larger than the agent, and the agent cannot represent the optimal value function and policy even in the limit of infinite data. In such scenarios, the agent needs to make decisions under significant uncertainty, and this presents several challenges, including exploration, generalization, and efficient learning. Even though the hypothesis suggests that incorporating side information might not be beneficial in learning the optimal policy and value in such scenarios, structural decomposition of large environments in different ways can allow benchmarking methods along different axes while allowing a deeper study into the performance of algorithms on parts of the environment that the agent has not yet experienced.
Modular decomposition can guide the agent’s exploration process by helping the agent explore different parts of the environment independently. Incorporating modularity opens up a gateway to novel methods and theoretical insights about the relationships between task decomposition, exploration, and learning efficiency in large environments. Relational decompositions can help the agent learn relationships between different entities, bolstering its ability to generalize to unseen parts of the environment. Finally, Structural information can be used to facilitate more efficient learning. For instance, in an auxiliary optimization pattern, the agent could learn faster by optimizing auxiliary tasks that are easier to learn or that provide useful information about the environment’s structure.
{mybox}
Patterns for Partial Observability and Big Worlds
  • Use temporal and state abstraction to abstract away temporal dependencies and non-Markovian aspects of the state. Utilize Modularity to tie these abstractions to learned primitives such as skills or options.
  • Use memory more efficiently as an abstraction or augmentation for learned transition models.
  • Warehouse policies and utilize them for optimization across timescales, such as in hierarchical methods, to make them more tractable.
  • Utilize modular decompositions for guiding separate and parallel exploration mechanisms for different parts of the state space. Utilize relational abstractions to make this knowledge more interpretable.
  • Utilize structure for task factorization to guide benchmarking methods along different axes of task complexity.
6.5 Automated RL
@@
Automated RL (AutoRL) is a sub-field focused on methods to automate the process of designing and optimizing RL algorithms, including the architecture of the agent, reward function, and other hyperparameters (?). Methods in AutoRL can be placed on a spectrum of automation, where on one end would be methods to select pipelines and on the other would be methods that try to discover new algorithms ground-up in a data-driven manner (?). Techniques from the Automated Machine Learning literature (?) then transfer to the RL setting, including algorithm selection (?), hyperparameter optimization (?, ?, ?), dynamic configurations (?), learned optimizers (?), and neural architecture search (?). Similarly, techniques from the Evolutionary optimization and Meta-Learning literature naturally transfer to this setting with methods aiming to meta-learn parts of the RL pipeline such as update rules (?), loss functions (?, ?), symbolic representations of algorithms (?, ?, ?), or concept drift (?). However, there are still many open questions in AutoRL, such as properties of hyperparameter landscapes in RL (?), sound evaluation protocols (?), stability of training due to the non-stationary learning task and non-deterministic data collection on the fly. Consequently, most of these methods suffer from a lack of scalability.
@@
Depending on the decomposability of the problem at hand, different RL methods could be more appropriate. Structural decompositions can guide the selection process in AutoRL by suggesting appropriate types of decompositions based on the problem characteristics. Understanding how different decomposition types influence the performance of RL methods can bridge the gap between selection and configuration by helping researchers understand the level of abstraction needed for selection conditioned on the task, aiding in the development of more efficient and targeted search algorithms. Decomposability can also guide ranking procedures, where methods that cater to different decomposability can be ranked differently, given a problem.
@@
Parameters related to structural decomposition (e.g., the number of subtasks in a modular decomposition) could be part of the hyperparameter optimization process in AutoRL. Researchers can investigate the interplay between configuration spaces of hyperparameters and various structural decomposition-related parameters. for example, high decomposability might require different exploration rates or learning rates than a low decomposability problem. This could lead to novel insights and methods for more effective hyperparameter optimization in AutoRL. Practitioners can use this understanding to guide the hyperparameter optimization process in their AutoRL system. By tuning parameters related to the decomposition, they can potentially improve the performance of their RL agent.
{mybox}
AutoRL Patterns
  • Use methods to perform a decomposability assessment of a problem. This can help guide algorithm selection for problems with different types of decomposability.
  • Expedite the hyperparameter search by abstracting away task-irrelevant aspects.
  • Warehouse to reuse learned policy and value functions for landmarking performances of algorithms similar to the one being optimized.
  • Incorporate modularity information in the form of goals and task hierarchies in the search process.
  • Structure neural architecture search-space using decomposability in the problem.
@@ 6.6 Meta-Reinforcement Learning
Meta-Reinforcement Learning, while having overlaps with AutoRL, is a field in and of itself (?) that focuses on training agents to quickly adapt and learn new tasks or environments. The general Meta-RL setup involves a bi-level optimization procedure where an agent learns a set of parameters by training on a distribution of tasks or environments that help it adapt and perform well on new, unseen tasks that share some form of overlap with the training tasks. ? (?) outline different problem settings in Meta-RL based on the kind of feedback (supervised, unsupervised, rewards) that is provided to the agent during the training and adaptation phases. We particularly refer to the standard setting where extrinsic rewards act as feedback during the training and adaptation phases. However, decompositions can also be useful for other settings similar to those discussed in other sections.
@@
Depending on the meta-task’s decomposability, different task decomposition approaches could be employed to guide the meta-learning process. Consequently, understanding how task decomposition affects Meta-RL can guide the development of more effective meta-learning algorithms. It might also lead to new insights on how to balance learning between different subtasks. By identifying suitable decompositions, practitioners can set up their system to learn in a way that is more aligned with the structure of the tasks, potentially leading to improved performance.
@@
Decompositions could inform the way a Meta-RL agent adapts to a new task. For instance, if the new task is highly decomposable, a modular adaptation strategy could be more appropriate by guiding the agent to an appropriate latent space of the new task. Thus, our framework of decomposability can inspire new research into how the task’s decomposition, f can guide adaptation strategies in Meta-RL. This could lead to novel methods or theories on adapting to new tasks more effectively based on their structure.
{mybox}
Meta-RL Patterns
  • Use decompositions for abstracting task distributions, which can then be integrated into the adaptation process.
  • Compartmentalize the learning process into modules for highly decomposable problems. These modules can serve as abstract configurations for the meta-level and, thus, make the outer loop more tractable
  • Learn and warehouse models geared towards specific task clusters for different decomposability types to guide data augmentation during the adaptation phase.
  • Utilize decomposability to design learning curricula based on abstract types of tasks to train the warmstarting configurations

### v3

@@ 7.2 Unsupervised RL
Structural decompositions can help such methods by improving the pre-training phase’s tractability and the fine-tuning phase’s generality. Latent decompositions could help exploit structure in unlabeled data, while relational decompositions could add interpretability to the learned representations. Through augmentation, conditioning policies on specific parts of the state space can reduce the data needed for fine-tuning. Additionally, understanding problem decomposition can simplify complex problems into more manageable sub-problems, effectively reducing the perceived problem complexity while incorporating such decomposition in external curricula for fine-tuning. Incorporating portfolios guided by decompositions for competence-based methods can boost the fine-tuning process of the learned skills.
Research Area 2: Structured Unsupervised RL Use latent decompositions to extract structure from unlabeled data, reducing Data Dependency. Employ factored and modular decompositions and abstractions to manage scalability by focusing learning on different parts of the problem independently. Store skills in a portfolio across different modular sub-problems to reuse solutions and enhance Generality. Manage Problem Complexity by leveraging problem decomposition to simplify the learning task and using decompositions for fine-tuning using curriculum learning.
@@ 7.3 Big Data and Foundation models in RL
Foundation models for RL come increasingly closer to becoming a reality. Such RL models would follow a similar concept of training a large model on various tasks, environments, and behaviors to be fine-tuned for specific downstream tasks. SMART (?), one of the current contenders for such models, follows this paradigm by using a self-supervised and control-centric objective that encourages the transformer-based model to capture control-relevant representation and demonstrates superior performance when used for fine-tuning. AdA (?) trains an in-context learning agent on a vast distribution of tasks where the task factors are generated from a latent ruleset.
Given the pre-training paradigm, these methods are highly data-dependent in principle. However, incorporating large amounts of data can demonstrate scalability benefits by reducing fine-tuning costs for distributed applications. A natural question is the role of Structured RL in the realm of end-end learning and big data. Even though such methods subscribe to an end-to-end paradigm, structural decompositions can benefit them differently.
Discovering Structure using Foundation Models
Foundation models offer avenues to discover structure in an unsupervised manner. Such methods have traditionally been prevalent in unsupervised RL, particularly by learning unsupervised skills. Integrating foundation models allows for learning predictive representations of different pipeline parts and discovering behaviors across modalities in different domains. Therefore, such methods can offer a good way to implicitly learn decompositions in problems and, consequently, utilize them using the aforementioned patterns.
@@
Researchers can better understand the decomposability in the pre-trained models by categorizing methods based on how they incorporate structure. Consequently, this can guide the selection processes for fine-tuning methods depending on the tasks. Passive learning from pre-trained models can benefit from better explanations about what parts of a fine-tuning task space might be suited for what kind of warmstarting strategies. Additionally, incorporating interpretability-oriented decompositions such as relational representations can help design more interpretable fine-tuning methods.
@@
With a better understanding of how different architectures and algorithms incorporate structural information, practitioners can more effectively adapt existing methods or contribute to designing novel solutions tailored to their specific tasks. For example, Action Branching architectures might provide modular functionality in downstream tasks, especially suited for multi-task settings. On the other hand, representation bottlenecks might suit settings that deviate from each other by small changes in contextual features.
@@
By understanding how to decompose tasks and incorporate structural information, foundation models can be fine-tuned more effectively for specific tasks. Understanding decompositions could guide how to structure the fine-tuning process or adapt the model to a new task.
@@
Problems with varying levels of decomposition might have specific benchmarking requirements. By accounting for these attributes, we can build better benchmarks and evaluation protocols for methods that utilize foundation models. Researchers can use this framework to design better evaluation protocols and benchmarks for foundational models. For practitioners, such benchmarks and evaluation protocols can guide the selection of models and algorithms for their specific tasks.
Research Area 3: Structure and Foundation Models for RL Use foundation models to discover decompositions in problems that can be subsequently used for fine-tuning and adaptation. Use Factored or Relational abstractions on the pre-trained foundation model for state abstractions to manage high-dimensional state spaces and, thus, reduce data dependency. Condition the policy on additional task-specific information, such as goal information, representation of specific fine-tuning instructions, or control priors to improve scalability. Regularize the fine-tuning process to prevent catastrophic forgetting of useful features learned during pre-training. Maintain a portfolio of fine-tuned policies and value functions to help reuse previously learned skills and adapt them to new tasks, improving learning efficiency and generalization. Incorporate a curriculum of increasingly complex fine-tuning environments based on the agent’s performance to help the agent gradually adapt the foundation model’s knowledge to the specific RL task. Use explicit architectures that fine-tune different RL problem aspects, such as perception, policy learning, and value estimation.
@@ 7.4 Partial Observability and Big Worlds
In many real-world situations, the Markov property might not fully capture the dynamics of the environment (?, ?). This can happen in cases where the environment’s state or the rewards depend on more than just the most recent state and action or if the agent cannot fully observe the state of the environment at each time step. In such situations, methods must deal with non-Markovian dynamics and partial observability.
@@
Abstractions can play a crucial role in such situations, where structural decompositions using abstraction patterns can make methods more sample-efficient. Often used in options, temporal Abstractions allow the agent to make decisions over extended periods, thereby encapsulating potential temporal dependencies within these extended actions. This can effectively convert a non-Markovian problem into a Markovian one at the level of options. State abstractions abstract away irrelevant aspects of the state and, thus, can sometimes ignore specific temporal dependencies, rendering the Markovian process at the level of the abstracted states. Thus, research into the role of decompositions in abstraction opens up possibilities to understand the dependencies between non-Markovian models and the abstractions they use to solve problems with incomplete information. Abstraction can also simplify the observation space in POMDPs, reducing the complexity of the belief update process. The abstraction might involve grouping similar observations, identifying higher-level features, or other simplifications. Abstractions allow us to break partial observability into different types instead of always assuming the worst-case scenario. Utilizing such restricted assumptions on partial observability can help us build more specific algorithms and derive convergence and optimality guarantees for such scenarios.
@@
Any additional information required, such as belief states or memories of past observations, can be used as abstractions or augmentations. This can also help with more efficient learning of transition models for planning. Hierarchical techniques that utilize optimization at different timescales can incorporate portfolios to reuse learned policies across various levels of abstraction. Environment generation patterns could also generate a curriculum of increasingly complex tasks for the agent, starting with simpler MDPs and gradually introducing partial observability or other non-Markovian features.
@@
As we extend the information content of the environment to its extremity, we delve into the realm of the big world hypothesis in RL (?), where the agent’s environment is multiple orders of magnitude larger than the agent. The agent cannot represent the optimal value function and policy even in the limit of infinite data. In such scenarios, the agent must make decisions under significant uncertainty, which presents several challenges, including exploration, generalization, and efficient learning. Even though the hypothesis suggests that incorporating side information might not be beneficial in learning the optimal policy and value in such scenarios, structural decomposition of large environments in different ways can allow benchmarking methods along different axes while allowing a deeper study into the performance of algorithms on parts of the environment that the agent has not yet experienced. Modular decomposition can guide the agent’s exploration process by helping the agent explore different parts of the environment independently. Incorporating modularity opens a gateway to novel methods and theoretical insights about the relationships between task decomposition, exploration, and learning efficiency in large environments. Relational decompositions can help the agent learn relationships between different entities, bolstering its ability to generalize to unseen parts of the environment. Finally, Structural information can be used to facilitate more efficient learning. For instance, in an auxiliary optimization pattern, the agent could learn faster by optimizing auxiliary tasks that are easier to learn or provide helpful information about the environment’s structure.
Research Area 4: Structure in Partial Observability and Big Worlds Use temporal and state abstraction to abstract away temporal dependencies and non-Markovian aspects of the state. Utilize Modularity to tie these abstractions to learned primitives such as skills or options. Use memory more efficiently as an abstraction or augmentation for learned transition models. Create a portfolio of policies and utilize them for optimization across timescales, such as in hierarchical methods, to make them more tractable. Utilize modular decompositions for guiding separate and parallel exploration mechanisms for different parts of the state space. Utilize relational abstractions to make this knowledge more interpretable. Utilize structure for task factorization to guide benchmarking methods along different axes of task complexity.
7.5 Automated Reinforcement Learning (AutoRL)
@@
Automated RL (AutoRL) is a sub-field focused on methods to automate the process of designing and optimizing RL algorithms, including the agent’s architecture, reward function, and other hyperparameters (?). Methods in AutoRL can be placed on a spectrum of automation, where on one end would be methods to select pipelines and on the other would be methods that try to discover new algorithms ground-up in a data-driven manner (?). Techniques from the Automated Machine Learning literature (?) then transfer to the RL setting, including algorithm selection (?), hyperparameter optimization (?, ?, ?), dynamic configurations (?), learned optimizers (?), and neural architecture search (?). Similarly, techniques from the Evolutionary optimization and Meta-Learning literature naturally transfer to this setting with methods aiming to meta-learn parts of the RL pipeline such as update rules (?), loss functions (?, ?), symbolic representations of algorithms (?, ?, ?), or concept drift (?). However, there are still many open questions in AutoRL, such as properties of hyperparameter landscapes in RL (?), sound evaluation protocols (?), stability of training due to the non-stationary learning task and non-deterministic data collection on the fly. Consequently, most of these methods suffer from a lack of scalability.
@@
Depending on the decomposability of the problem at hand, different RL methods could be more appropriate. Structural decompositions can guide the selection process in AutoRL by suggesting appropriate types of decompositions based on the problem characteristics. Understanding how different decomposition types influence the performance of RL methods can bridge the gap between selection and configuration by helping researchers understand the level of abstraction needed for selection conditioned on the task, aiding in developing more efficient and targeted search algorithms. Decomposability can also guide ranking procedures, where methods that cater to different decomposability can be ranked differently, given a problem.
@@
Parameters related to structural decomposition (e.g., the number of subtasks in a modular decomposition) could be part of the hyperparameter optimization process in AutoRL. Researchers can investigate the interplay between configuration spaces of hyperparameters and various structural decomposition-related parameters. For example, a high decomposability problem might require different exploration or learning rates than a low one. This could lead to novel insights and methods for more effective hyperparameter optimization in AutoRL. Practitioners can use this understanding to guide the hyperparameter optimization process in their AutoRL system. By tuning parameters related to the decomposition, they can potentially improve the performance of their RL agent.
Research Area 5: Structure in AutoRL Use methods to perform a decomposability assessment of a problem. This can help guide algorithm selection for problems with different types of decomposability. Expedite the hyperparameter search by abstracting away task-irrelevant aspects. reuse learned policy and value functions stored in a portfolio for landmarking performances of algorithms similar to the one being optimized. Incorporate modularity information in the form of goals and task hierarchies in the search process. Structure neural architecture search-space using decomposability in the problem.
@@ 7.6 Meta-Reinforcement Learning
Meta-reinforcement learning, while having overlaps with AutoRL, is a field in and of itself (?) that focuses on training agents to adapt and learn new tasks or environments quickly. The general Meta-RL setup involves a bi-level optimization procedure where an agent learns a set of parameters by training on a distribution of tasks or environments that help it adapt and perform well on new, unseen tasks that share some form of overlap with the training tasks. ? (?) outline different problem settings in Meta-RL based on the kind of feedback (supervised, unsupervised, rewards) provided to the agent during the training and adaptation phases. We mainly refer to the standard setting where extrinsic rewards act as feedback during the training and adaptation phases. However, decompositions can also be helpful for settings like those discussed in other sections.
@@
Different task decomposition approaches could guide the meta-learning process depending on the meta-task’s decomposability. Consequently, understanding how task decomposition affects Meta-RL can guide the development of more effective meta-learning algorithms. It might also lead to new insights on balancing learning between different subtasks. By identifying suitable decompositions, practitioners can set up their system to learn in a way that is more aligned with the structure of the tasks, potentially leading to improved performance.
@@
Decompositions could inform the way a Meta-RL agent adapts to a new task. For instance, if the new task is highly decomposable, a modular adaptation strategy could be more appropriate by guiding the agent to an appropriate latent space of the new task. Thus, incorporating existing patterns can inspire new research into how the task’s decomposition can guide adaptation strategies in Meta-RL. This could lead to novel methods or theories on adapting to new tasks more effectively based on their structure.
Research Area 6: Structured Meta-RL Use decompositions for abstracting task distributions, which can be integrated into the adaptation process. Compartmentalize the learning process into modules for highly decomposable problems. These modules can serve as abstract configurations for the meta-level and, thus, make the outer loop more tractable Learn and create a portfolio of models geared towards specific task clusters for different decomposability types to guide data augmentation during the adaptation phase. Utilize decomposability to design learning curricula based on abstract types of tasks to train the warmstarting configurations